OpenStudy (anonymous):

Solve. Explain. x+ y = 52 x-y = 16 (Systems of equations --Substitution method)

OpenStudy (anonymous):

here what you want to do is subtract the two equations from each other to eliminate one of the variables

OpenStudy (anonymous):

so X is already set up for you to minus them out

OpenStudy (anonymous):

Could you visually solve this problem too..after you're done enplaning it..? I'm more of a visual learner..

OpenStudy (anonymous):

X-X=0, Y-(-Y)=2Y, and 52-16=36

OpenStudy (anonymous):

ya sure

OpenStudy (anonymous):

now you have 2Y=36

OpenStudy (anonymous):

Y=18

OpenStudy (anonymous):

plug 18 in for Y in the original equation and solve for X

OpenStudy (anonymous):

X+18=52

OpenStudy (anonymous):

I would have added so that y is eliminated. You make less stupid mistakes when you add.

OpenStudy (anonymous):

But it doesn't really matter

OpenStudy (anonymous):

X=34

OpenStudy (anonymous):

I like to get rid of the left most term, habit I guess

OpenStudy (anonymous):

x+y=52 x-y=16 ------- 2x=68 or,x=68/2 x=34

OpenStudy (anonymous):

then the value of x in 1 equation x+y=52 or,34+y=52 or.y=52-34 y=18 ans